OpenStudy (anonymous):

What is the factored form of the expression? 3x^2 + 8x – 16

sam (.sam.):

For a polynomial of the form ax^(2)+bx+c, find two factors of a*c (-48) that add up to b (8).In this problem 4*-(4)/(3)=-(16)/(3) (which is (c)/(a)) and 4-(4)/(3)=(8)/(3) (which is ((b)/(a)) , so insert 4 as the right hand term of one factor and -(4)/(3) as the right-hand term of the other factor. (x+4)(x-(4)/(3))

sam (.sam.):

That's much more clearer

OpenStudy (apoorvk):

3x^2 +12x - 4x - 16 = 3x(x+4) - 4(x+4) = (3x-4)(x+4).